In a pilot study with 20 subjects evaluating the use of a new drug to lower...

  1. In a pilot study with 20 subjects evaluating the use of a new drug to lower resting heart rates (HR) of patients, the following data was recorded:

Subject #

Resting HR

001

72

002

88

003

71

004

87

005

64

006

77

007

79

008

59

009

77

010

68

011

78

012

89

013

91

014

81

015

85

016

75

017

69

018

75

019

77

020

81

Given that the average resting HR of the general population for this study is 72.  perform the appropriate t test. Use the two-tailed test. Explain why the textbook tells us to start with the two-tailed test. What is the value of t? Using an alpha of 0.05, is the t statistic significant? Why? What are the confidence limits for a 95% confidence interval here and what do they mean for this patient group

Now that you have answered question 6 with the two-tailed test, answer it with the one-tailed test.

How does this change your answers to our conventional questions:   What is the value of t? What is the p value? Why did this change? Using an alpha of 0.05, is the t statistic significant? Why?
Answer #1

ar x = (72 + 88 + 71 + 87 + 64 + 77 + 79 + 59 + 77 + 68 + 78 + 89 + 91 + 81 + 85 + 75 + 69 + 75 + 77 + 81)/20 = 77.15

s = sqrt(((72 - 77.15)^2 + (88 - 77.15)^2 + (71 - 77.15)^2 +( 87 - 77.15)^2 + (64 - 77.15)^2 + (77 - 77.15)^2 + (79 - 77.15)^2 + (59 - 77.15)^2 + (77 - 77.15)^2 + (68 - 77.15)^2 + (78 - 77.15)^2 + (89 - 77.15)^2 + (91 - 77.15)^2 + (81 - 77.15)^2 + (85 - 77.15)^2 + (75 - 77.15)^2 + (69 - 77.15)^2 + (75 - 77.15)^2 + (77 - 77.15)^2 + (81 - 77.15)^2)/19) = 8.5

For two-tailed test

H0: mu = 72

H1: mu eq 72

The test statistic t = (ar x - mu)/(s/sqrt n)

                             = (77.15 - 72)/(8.5/V20)

                             = 2.71

df = 20 - 1 = 19

P-value = 2 * P(T > 2.71)

              = 2 * (1 - P(T < 2.71))

              = 2 * (1 - 0.9931)

              = 2 * 0.0069

              = 0.0138

At alpha = 0.05, since the P-value is less than alpha(0.0138 < 0.05), so we should reject the null hypothesis.

For one-tailed test

H0: mu = 72

H1: mu < 72

The test statistic t = (ar x - mu)/(s/sqrt n)

                             = (77.15 - 72)/(8.5/V20)

                             = 2.71

df = 20 - 1 = 19

P-value = P(T > 2.71)

              = 1 - P(T < 2.71)

              = 1 - 0.9931

              = 0.0069

At alpha = 0.05, since the P-value is less than alpha(0.0069 < 0.05), so we should reject the null hypothesis.